# Automated Layout Design and Control of Robust Cooperative Grasped-Load Aerial Transportation Systems

### v2

@@ II-B Layout Optimization
Figure 4: Control infrastructure diagram
II-C Control Infrastructure
@@
The dynamics of motion of each quadcopter is highly constrained due to the rigid attachments. Therefore, using a traditional quadcopter controller would be challenging. However the integrated system corresponds itself to a larger rigid body (neglecting effects due to internal elasticity and vibrations). Therefore, approaching the overall dynamics and control as a whole (as was done for the placement optimization) is preferable. This strategy leads to a hierarchical control infrastructure, in which a single estimator is used for the overall motion and the quadcopters receive thrust commands individually through $N$ parallel communication channels. Furthermore, this control strategy comes naturally from the layout optimization computation, during which the optimal LQR gains have been computed. A block diagram of the suggested control infrastructure is shown in Fig. 4. The overall behavior of the controller is dominated by the choice of some parameters which directly impact the computation of the $\mathcal{H}_{2}$ cost. First, the disturbance matrix $B_{d}$ (eq. 1) describing how the random disturbance affects the dynamics, and, in particular, its variance along different directions in the wrench space. Second, the matrices $C$ and $D$ (eq. 2), i.e. how the auxiliary variable $z$ is defined. For simplicity, the $B_{d}$ matrix was chosen to weigh the six disturbance components equally, i.e. $B_{d}(4,1)=B_{d}(5,2)=B_{d}(6,3)=B_{d}(10,4)=B_{d}(11,5)=B_{d}(12,6)=1$. The $C$ matrix is chosen to just weight the position of the payload centroid and the system’s yaw angle (i.e. $C\in\mathbb{R}^{(4+4N)\times 12}$). After evaluating the system’s behavior in simulation, we set $C(1,1)=C(2,2)=0.5$ (corresponding to the position on a horizontal plane $x-y$), $C(3,3)=10$ (corresponding to vertical position $z$), $C(4,9)=50$ (corresponding to yaw angle), and zero for all the other components. The input weight matrix $D\in\mathbb{R}^{(4+4N)\times 4N}$ was set to the block matrix $D=[\mathbf{0}_{4N\times 4},I_{4N}]^{T}$, where $\mathbf{0}_{m\times n}$ denotes the $m\times n$ matrix of all zero elements and $I_{k}$ the $k$-dimensional identity.
II-D Hardware Setup
The vehicles we used as a reference for layout design and eventually for experiments are quadcopters equipped with a PX-4 flight controller. The relevant vehicle parameters can be found in Tab. I.
@@ II-D Hardware Setup
TABLE I: Table of parameters
The battery capacity ($5300\,\mathrm{mAh}$) has been chosen to guarantee $10$ minutes of flight time at maximum payload. The system theoretically is able to lift an external payload of $\sim 0.8\,\mathrm{kg}$ per quadcopter. It is interesting to note that the main contribution to the overall inertia of the system is due to the quadcopters, since they are attached to the payload perimeter, and thus at a larger distance from the overall center of mass. It is interesting to highlight that contributing a greater inertia than the actual payload transported is in line with all the ground and aerial transportation systems in use nowadays.
III RESULTS
The optimization framework which we developed allows us to determine the optimal placement of quadcopters for transportation across many different payload shapes, mass values and number of quadcopters used. In Fig. 5 we show some examples of optimal placements using the vehicles described in Tab. I as a reference. It is interesting to observe that the quadcopters are placed in a way which seems to trade off two different contributions. On the one hand, the area of the polygon defined by the quadcopters’ attachment points tends to be maximized in order to maximize control authority. On the other hand the distances of the attachment points to the system’s center of mass tend to be equalized. This is an effect of employing the Mahalanobis distance as a metric. In fact, even if the control authority is maximized (i.e. the input covariance $\Sigma_{\mathbf{u}}$ has smaller eigenvalues), we also need the quadcopters to have a comparable margin between feedforward thrusts $\bar{\mathbf{u}}$ and bounds, otherwise the Mahalanobis distance objective is decreased.
@@ III RESULTS
Figure 5: Optimal layouts for various object shapes and masses using different number of quadcopters. Our approach can handle both convex and non-convex polygons. It is interesting to note that, in contrast to common intuition, the optimal drone placement around a symmetric shape might not be symmetric as well.
IV EXPERIMENTAL VALIDATION
@@
For the flight experiments the quadcopters were attached to two different square panels of different masses and sizes. A heavier panel (A) of side length $0.61\,\mathrm{m}$ and mass $2.54\,\mathrm{kg}$ was used to test the payload capacity, and a lighter one (B) of side length $0.45\,\mathrm{m}$ and mass $1.088\,\mathrm{kg}$ to test the disturbance rejection performances. The rigid attachments have been obtained by screwing directly the quadcopters to the panels through wood dowels. The experiments are conducted in an indoor flight space equipped with a Motion Capture system, and markers for motion tracking are placed directly on the panels.
IV-A Payload lifting
Panel A was used to test the payload capacity of the system. The four quadcopters were attached to the corners of the panel and a hovering position was commanded. The system was able to successfully take off, stabilize, hover and land safely. An important observation to highlight is the close proximity of the feedforward thrusts $\bar{\mathbf{u}}$ to the upper thrust limit $\mathbf{u}_{h}$. Given the significant mass of Panel A ($\sim 2.5\,\mathrm{kg}$) compared to the quadcopter masses ($\sim 0.95\,\mathrm{kg}$ including battery), the narrow margin causes the system to have limited disturbance response capabilities, especially within the spatial limits of the flight space available for experiments.
IV-B Disturbance Rejection
@@
The four quadcopters were attached to panel B using two distinct configurations: an optimal one, placed at the corners, and a sub-optimal one. These configurations can be seen in Fig. 6. As a disturbance, a weight of $0.5,\mathrm{kg}$, or approximately $10\%$ of the system’s total mass, was attached to a corner of the panel during the flight, specifically under Quadcopter 1 (Fig. 6). The mass attachment is done through a magnet to guarantee the repeatability of the experiment. Both the state and input data were collected and compared to validate their consistency with the layout optimization framework. Representative response plots over time are presented in Fig. 7.
@@
From the data, it is clear that the choice of quadcopter layout has a significant impact on disturbance rejection performance. The optimal layout results in lower oscillations in the attitude of the vehicle, specifically in the pitch and roll angles, and also recovers more quickly from disturbances. In comparison, the sub-optimal layout sees its peak deviation from hovering conditions nearly double the one of the optimal setup. Another observation is related to thrust behavior after adding the payload. In the sub-optimal configuration, the increase in thrust following the addition of the payload is more noticeable. Additionally, the steady-state thrust from the quadcopter positioned closest to the added payload is also higher in the sub-optimal setup. This suggests that, despite facing disturbances of the same magnitude, the thrust distribution becomes more uneven in the sub-optimal case, emphasizing the benefits of an optimal quadcopter layout. It is important to note that in the sub-optimal configuration, the model and controller gains were recomputed from scratch to maximize performances (i.e. we did not use the same controller as in the optimal configuration).
Figure 6: Attachment layouts tested on panel B for disturbance rejection comparison. Left: optimal configuration, right: sub-optimal configuration.
V CONCLUSIONS
In this study, we introduced a novel approach to cooperative aerial transportation with a team of quadcopters, employing optimal control theory and a hierarchical control strategy. This methodology offers an original approach in which the layout design and control problems are solved in a coupled way through an optimization routine derived from $\mathcal{H}_{2}$ control theory.
Figure 7: Response to step disturbance over time. The mass is attached instantly (through a magnet) underneath Quadcopter 1. The attachment time is indicated with a vertical dashed line. On the top row, attitude response. On the bottom row, total quadcopter thrusts. Left: optimal configuration, right: sub-optimal configuration
Our optimization framework allows to cater our system to different payload shapes, inertias, and quadcopter counts. The hierarchical control structure offers a systematic and simple way to deal with the dynamics viewing the system as a single, integrated rigid body. The choice of the weight matrices $C$ and $D$ represents the freedom given to the control designer to obtained the desired performances. Overall, the main contributions of this paper are:
  • An automated tool to optimally arrange thrust modules on a payload to maximize robustness of the integrated system;
  • A control framework general enough to be able to fly different payload shapes and masses using the preferred number of quadcopters.
We also emphasize the idea that dealing with system design and control in a coupled manner simplifies the control complexity and increases the overall system performances compared to a sequential siloed approach (control design following system design).
@@ V CONCLUSIONS
Experimental results validate the effectiveness of our approach. Flight tests under disturbance highlighted the stability and robustness of the system, and, most importantly, the agreement between the performance predicted by our optimization tool and experimental evidence.
Future developments for this work include:
  • Exploring more advanced control strategies and infrastructures. At this stage, for example, the vehicles’ onboard sensor data were not used.
  • Removing the assumption of known payload inertia and developing a control scheme under this uncertainty component.
  • Using the system as a moving tray that can be used to transport and balance a payload placed on its top.
In conclusion, this study highlights the coupling between design variables and control. The physical placement of quadcopters significantly affect the performances of the controller. Optimizing the layout allows to take more advantage of optimal control tools rather than applying them at a second stage, once the design process is finished.
VI ACKNOWLEDGMENTS
This work was funded by the Tsinghua-Berkeley-Shenzhen Institute (TBSI). The authors would also like to thank Roman Ibrahimov and Ruiqi Zhang for their precious help.

### v3

@@ II-B Control and layout co-design
II-C Control Implementation
For simplicity, we treat the integrated quadcopter-payload system as a single rigid body, and use the LQR gains computed in the co-design optimization algorithm 1, which are readily available as a byproduct of our co-design tool. This approach leads to a hierarchical control infrastructure, in which a single estimator is used for the overall motion and the quadcopters receive thrust commands individually through $N$ parallel communication channels. A block diagram of the control infrastructure is shown in Fig. 4.
An alternative control infrastructure, allowing compatibility with off-the-shelf autopilots, could be commanding a specific total thrust for each quadcopter. This approach would however be strictly suboptimal compared to independently commanding each motor. More advanced control strategies retaining the same first-order properties achieved by the $\mathcal{H}_{2}$ controller could also be used. One example could be a Model Predictive Control approach using the same optimal LQR weights as our co-design tool. Using alternative controllers is beyond the co-design focus of this work.
@@ II-C Control Implementation
The overall behavior of the controller is dominated by the choice of some hyperparameters which directly impact the computation of the $\mathcal{H}_{2}$ cost. First, the disturbance matrix $B_{d}$ from (2), that describes how the random disturbance $\mathbf{d}$ affects the dynamics, and, in particular, quantifies its covariance along different directions in the wrench space. Second, the matrices $C$ and $D$ from (3), i.e. how the auxiliary variable $z$ is defined. For simplicity, the $B_{d}$ matrix was chosen to weigh the six disturbance components equally, i.e. $B_{d}(4,1)=B_{d}(5,2)=B_{d}(6,3)=B_{d}(10,4)=B_{d}(11,5)=B_{d}(12,6)=1$. The $C$ matrix is chosen to just weigh the position of the payload centroid and the system’s yaw angle (i.e. $C\in\mathbb{R}^{(4+4N)\times 12}$). After evaluating the system’s behavior in simulation, we set $C(1,1)=C(2,2)=0.5$ (corresponding to the position on a horizontal plane $x-y$), $C(3,3)=10$ (corresponding to vertical position $z$), $C(4,9)=50$ (corresponding to yaw angle), and zero for all the other components. The input weight matrix $D\in\mathbb{R}^{(4+4N)\times 4N}$ was set to the block matrix $D=[\mathbf{0}_{4N\times 4},I_{4N}]^{T}$, where $\mathbf{0}_{m\times n}$ denotes the $m\times n$ matrix of all zero elements and $I_{k}$ the $k$-dimensional identity.
II-D Optimization Results
The optimization framework we developed allows us to determine the optimal placement of quadcopters for transportation across different payload shapes, mass values and number of quadcopters used. In Fig. 5 we show some examples of optimal placements using the vehicles described in Tab. I as a reference. It is interesting to observe that the quadcopters are placed in a way which seems to trade off two different contributions. On the one hand, the area of the polygon defined by the quadcopters’ attachment points tends to be maximized in order to maximize control authority. On the other hand the distances of the attachment points to the system’s center of mass tend to be equalized. This is an effect of employing the Mahalanobis distance as a metric. In fact, even if the control authority is maximized (i.e. the input covariance $\Sigma_{\mathbf{u}}$ has smaller eigenvalues), we also need the quadcopters to have a comparable margin between feedforward thrusts $\bar{\mathbf{u}}$ and saturation, otherwise the Mahalanobis distance objective is decreased. Another interesting observation is that a symmetric object does not lead to a symmetric optimal configuration necessarily, and sometimes the optimum can be counterintuitive. This is shown by the concave shape in Fig. 5.
Figure 4: Diagram of the control infrastructure. The specific loop rates are relative to the available experimental testbed.
TABLE I: Table of quadcopter parameters
@@ II-D Optimization Results
Figure 5: Optimal layouts for various object shapes and masses using different number of quadcopters. Our approach can handle both convex and non-convex polygons. It is interesting to note that, in contrast to common intuition, the optimal drone placement around a symmetric shape might not be symmetric.
III Experiments
The relevant parameters of the vehicles used are shown in Tab. I. The battery capacity ($5300\,\mathrm{mAh}$) has been chosen to guarantee $10$ minutes of flight time at maximum payload. The system theoretically is able to lift an external payload of up to $\sim 0.8\,\mathrm{kg}$ per quadcopter. It is interesting to note that the main contribution to the overall inertia of the system is due to the quadcopters, since they are attached to the payload perimeter, and thus at a larger distance from the overall center of mass. It is also important to highlight that contributing a greater inertia than the actual payload is in line with all the ground and aerial transportation systems in use nowadays. In fact, most of the transportation systems that we use daily (e.g. trains, planes, cars) have significantly greater mass and inertia compared to the payload they transport.
For the flight experiments three different payload shapes were used : a square panel (A) of mass $1.02\,\mathrm{kg}$, a concave square (B) of mass $0.71\,\mathrm{kg}$, and an L-shaped panel of mass $0.76\,\mathrm{kg}$, all of side length $0.45\,\mathrm{m}$. For these three options, the performance of optimal quadcopter configurations were tested against some suboptimal case (see Fig. 6). In the case of Panel B, a symmetric layout was chosen as the suboptimal configuration. Panel A was lifted by four quadcopters, while for panel B and C only three were used. On panels A and B, we performed hovering tests (with and without wind), response to position reference step tests, and disturbance rejection tests when suddenly attaching a disturbance mass. Panel C was used for a trajectory tracking experiment. Our experiments aim at showing the comparative advantages achieved through the proposed method, rather than focusing on the absolute performance metrics of the controller, which can be optimized through various other state-of-the-art methods.
Figure 6: Attachment layouts tested on panel A, B, and C. Top row: optimal configurations, bottom row: suboptimal configuration. The red cross indicates where the disturbance mass is attached in the disturbance rejection experiments on A and B. Panel C is used for trajectory tracking experiments.
@@ III Experiments
The rigid attachments were obtained by directly bolting the quadcopters to the panels through wood dowels. The experiments are conducted in an indoor flight space equipped with a Motion Capture system, and markers for motion tracking are placed directly on the panels. A picture of the system flying in the flight space is shown in Fig. 1.
The square panel (A) was also used to test the payload capacity of the system. Four quadcopters attached at the corners were able to lift, hover and land with up to $\sim 2.5\,\mathrm{kg}$. Large payload masses, however, are not ideal due to the narrow margin from actuator saturations. The reduced control authority causes the system to have limited disturbance response capabilities, especially within the spatial limits of the flight space available for experiments. Supplementary videos showing relevant experiments can be found at https://hiperlab.berkeley.edu/members/carlo-bosio/automatedlayout/.
III-A Hovering
As a first comparison, we report the hovering performances of the different systems in the various configurations we tested. The hovering was performed in still air and with a $1\,\mathrm{m\cdot s^{-1}}$ wind disturbance. The RMSE values, averaged across five runs roughly $15$ seconds long, of position deviation from reference and attitude angles can be found in Tabs. II and III. The optimal configurations for both panels are more stable and less subject to disturbance. The difference is particularly significant for Panel (B), which could not withstand the wind disturbance at all, and showed an unstable behavior (due to actuator saturation) in all the five trials.
Real world systems are inevitably affected by noise (e.g., imperfect sensors, communication delays, small modeling inaccuracies). Additionally, due to the sizes of the system ($\sim 1.5\,\mathrm{m}$) compared to the size of the flight space ($\sim 5\,\mathrm{m}$), significant undesired aerodynamic perturbations are present. These disturbances are unavoidable and partially explain the imperfect control absolute performances. However, they are present in the same amount regardless of the specific drone configuration. The comparative analysis suggests that the configuration plays an important role in improving the disturbance robustness.
We also compare the hovering performances of an $\mathcal{H}_{\infty}$ controller on the optimal configuration for Panel (B) (Fig. 6). We simulate the system hovering subject to the nominal Gaussian noise for 10 minutes (for convergence of the RMSE statistics). The RMSE values are at least $50\%$ lower for the nominal $\mathcal{H}_{2}$ controller. We experimentally verify this by testing the system, controlled through $\mathcal{H}_{\infty}$, hovering in presence of $1\mathrm{m\cdot s^{-1}}$ wind. The experiments confirm the higher performances of the $\mathcal{H}_{2}$ control with similar differences in the RMSE values. This is expected as the worst-case disturbances considered by $\mathcal{H}_{\infty}$ synthesis are likely not observed in this setting. In a scenario like ours, where the disturbance is reasonably modeled as white noise, the $\mathcal{H}_{2}$ approach leads to higher performance. Arguments and results of similar nature can be found in the literature [37, 38].
No wind
Panel (A) (B)
[…]
$x\,[\mathrm{m}]$ 0.284 0.1390 51 0.320 0.0547 83
$y\,[\mathrm{m}]$ 0.1440 0.0775 46 0.1240 0.0909 27
$z\,[\mathrm{m}]$ 0.0713 0.0674 5.5 0.0690 0.0619 10
$Y\,[\mathrm{rad}]$ 0.0547 0.0537 1.8 0.0687 0.0608 12
$P\,[\mathrm{rad}]$ 0.0235 0.0163 31 0.0231 0.0146 37
$R\,[\mathrm{rad}]$ 0.0213 0.0149 30 0.0206 0.0161 21
TABLE II: Table of RMSE values for position and attitude averaged across five different undisturbed hovering trials for panels (A) and (B), with the respective suboptimal (S-opt. column) and optimal (Opt. column) configurations. The relative improvement of performances can be seen in the bold columns, in percentage.
$1\,\mathrm{m\cdot s^{-1}}$ wind
Panel (A) (B)
[…]
$x\,[\mathrm{m}]$ 0.6374 0.216 66 N/A 0.1285 -
$y\,[\mathrm{m}]$ 0.1520 0.1373 9.6 N/A 0.1144 -
$z\,[\mathrm{m}]$ 0.5203 0.465 10.5 N/A 0.0773 -
$Y\,[\mathrm{rad}]$ 0.1168 0.0890 24 N/A 0.1031 -
$P\,[\mathrm{rad}]$ 0.0257 0.0182 29 N/A 0.0282 -
$R\,[\mathrm{rad}]$ 0.0232 0.0156 33 N/A 0.0407 -
TABLE III: Table of RMSE values for position and attitude averaged across five different hovering trials when a $1\,\mathrm{m\cdot s^{-1}}$ wind is acting on the system. Data are reported for panels (A) and (B), with the respective suboptimal (S-opt. column) and optimal (Opt. column) configurations. The relative improvement of performances can be seen in the bold columns, in percentage. The suboptimal configuration for panel (B) was not able to withstand the wind disturbance, due to thrust saturations.
III-B Reference Step
As a second performance test, with the system hovering, a reference step change of $1\,\mathrm{m}$ is applied, and the transient response is recorded. In Fig. 7 the response of the square Panel (A) system to an applied position reference step (in particular a $1\,\mathrm{m}$ offset along the $x$ direction) is shown. It can be observed that the optimal configuration has a lower overshoot, but also a faster transient. After six seconds, the suboptimal configuration has not yet fully recovered from the event.
III-C Trajectory tracking
We also conducted trajectory tracking experiments on the L-shaped panel (C in Fig. 6). The commanded trajectory was a circle of diameter $2\,\mathrm{m}$ within the horizontal plane of height $z=2\,\mathrm{m}$. The reference velocity is $0.5\,\mathrm{m\cdot s^{-1}}$. Some example trajectories obtained are shown in Fig. 8, which show substantial qualitative differences in the tracking performances.
III-D Disturbance Rejection
@@
As a third performance test, a disturbance mass was attached underneath the panels during flight, specifically at the attachment points shown by the red cross signs in Fig. 6. The mass attachment is done through a magnet to guarantee the repeatability of the experiment across different runs. Both the state and input data were collected and compared to validate their consistency with the layout optimization framework. For Panel (A), the mass was of about $0.5\,\mathrm{kg}$ (approximately $10\%$ of the total mass), whilst for Panel (B), it was of about $0.3\,\mathrm{kg}$ ($7\%$ of the total mass). Once again, in the case of Panel (B), after a significant transient, most of the time the suboptimal configuration was not able to re-stabilize after the application of the disturbance.
@@
Representative response plots over time are presented in Fig. 9. From the data, it is clear that the choice of quadcopter layout has a significant impact on disturbance rejection performance. The optimal layout leads to lower oscillations in the attitude of the vehicle, specifically in the pitch and roll angles, and also recovers more quickly from disturbances. In comparison, the suboptimal layout sees its peak deviation from hovering conditions nearly double the one of the optimal setup. Another observation is related to thrust behavior after adding the payload. In the suboptimal configuration, the increase in thrust following the addition of the payload is more noticeable. Additionally, the steady-state thrust from the quadcopter positioned closest to the added payload is also higher in the suboptimal setup. This suggests that, despite facing disturbances of the same magnitude, the thrust distribution becomes more uneven in the suboptimal case, emphasizing the benefits of an optimal quadcopter layout. Finally, the bottom row of Fig. 9 shows the minimum Mahalanobis distance (defined in (10) and depicted in Fig. 3) from actuator saturations. The optimal layouts, in general, maintain a greater distance from this unsafe condition.
Figure 7: $x-z$ positions of the optimal and suboptimal configurations on Panel (A). The position reference step consists of an offset of $1\,\mathrm{m}$ along $x$. The optimal configuration has lower overshoot and faster recovery.
IV Conclusion
In this study, we introduced a novel approach to the layout and control co-design of cooperative aerial transportation systems. This methodology jointly solves the layout design and control synthesis problems through an optimization routine inspired from $\mathcal{H}_{2}$ control theory. Our optimization framework allows to cater our system to different payload shapes, inertias, and quadcopter counts. The choice of the weight matrices $C$ and $D$ represents the freedom given to the control designer to obtain the desired performances. Overall, the main contribution of this paper is a co-design tool using ideas from robust control to compute the optimal layout of thrust modules around a payload. The two main components of our approach are:
  • A novel cost function able to capture the robustness to disturbances of the multi-UAV system;
  • An efficient procedure to optimize this cost function with respect to the layout variables.
We also emphasize the idea that dealing with physical parameters and control jointly increases the overall system performances compared to a sequential, siloed approach (control design following layout design), sometimes even making the difference between failure and success. Moreover, extending this to nonlinear controllers could be done as follows: a desired nonlinear control strategy is implemented, and the control gains are selected so that the linearized feedback matches the computed gains. Of course, arguments about saturation become weaker, as saturation is likely to occur in areas where linearization is no longer accurate.
@@ IV Conclusion
The experimental results validate the effectiveness of our approach, even when fine engineering effort are not made to estimate system parameters and develop sophisticated controllers. Flight tests under disturbance highlighted the robustness of the system, and the agreement between the performance predicted by our optimization tool and experimental evidence.
Possible extensions for this work include exploring more advanced control strategies for the deployment stage and doing online payload parameter estimation in cases of uncertainty.
Figure 8: Example trajectory tracking experiments using the L-shaped panel (C). The commanded trajectory is a circle of diameter $2\,\mathrm{m}$ within the horizontal plane of height $z=2\,\mathrm{m}$.
To conclude, our study showed that the co-design of control and physical parameters in robotics is a promising path to achieve superior performance and reliability.
Figure 9: Response to step disturbance (added mass) over time. The mass is attached through a magnet, respectively where indicated by the red crosses in Fig. 6. The attachment instant is indicated with a vertical dashed line in the plots. On the top row: position deviation from set point. On the mid row: attitude response. On the bottom row: mean thrusts across each quadcopter (with dispersion bands in grey). Left: optimal and suboptimal configurations for panel (A), right: optimal and suboptimal configurations for panel (B). The grey area labeled as ‘cnt’ in the bottom row represents the contours of the individual propeller thrusts (not averaged across each quadcopter). As it is possible to observe, the suboptimal configurations have wider contours. The bottom row shows the minimum Mahalanobis distance (defined in (10) and depicted in Fig. 3) from the saturations throughout the experiments (higher is better). The optimal placement of the quadcopters allows the system to overall maintain a greater normalized distance from the actuator saturations.
Acknowledgments
This work was supported by the Tsinghua-Berkeley-Shenzhen Institute (TBSI), and the Powley fund of the Mechanical Engineering department of UC Berkeley. The experimental testbed at the HiPeRLab is the result of contributions of many people, a full list of which can be found at hiperlab.berkeley.edu/members/.